# A Cognitive Writing Perspective for Constrained Long-Form Text Generation

### v2

@@ A Cognitive Writing Perspective for Constrained Long-Form
Kaiyang Wan1, Honglin Mu1, Rui Hao2, Haoran Luo3, Tianle Gu1, Xiuying Chen1thanks: Corresponding author., 1Mohamed bin Zayed University of Artificial Intelligence, 2University of Chinese Academy of Sciences, 3Nanyang Technological University {Kaiyang.Wan, Xiuying.Chen}@mbzuai.ac.ae
@@ 1 Introduction
LLMs like ChatGPT Achiam et al. (2023) have begun to mirror human-like writing capabilities across diverse natural language processing tasks Xi et al. (2023). From crafting concise summaries Zhang et al. (2024b) to composing structured reports Schmidgall et al. (2025); Wang et al. (2024d), these models can generate coherent text in a single pass Rasheed et al. (2025); Minaee et al. (2024) with a fluency that often rivals human writers. Recent advances have led to models with expanded context windows of up to 128K tokens Pawar et al. (2024), theoretically enabling the generation of extensive documents Bai et al. (2024). However, these models face significant challenges when tasked with generating constrained long-form text under complex constraints, such as following detailed instructions over 10,000 words Wu et al. (2024a). This limitation poses a crucial barrier for applications requiring extended Shi et al. (2024), well-structured content, including creative design proposals, technical documentation, and comprehensive research reports.
@@ 2 A Cognitive Writing Perspective
However, LLMs lack this ability due to their left-to-right, single-pass generation Yao et al. (2023); Wu et al. (2024b). Without the ability to revisit or reorganize previous content, LLMs struggle with global revisions, such as restructuring sections or ensuring consistency across distant parts of the text Bae and Kim (2024); Cheng et al. (2024); Zhang et al. (2024a). This absence of dynamic reviewing often results in long-form outputs with accumulated errors, inconsistencies, or redundant content.
@@ 4 Methodology
Drawing upon our analysis of cognitive writing processes and the identified limitations of single-pass generation approaches, in this section, we propose CogWriter, a training-free framework that equips LLM with cognitive writing capabilities.
@@ 5.1 Experimental Setup
Models
We first present the performance of several single-pass generation baseline models from the official LongGenBench repository, including LongWriter-Llama3.1-8B Bai et al. (2024), Llama-3.1-8B-Instruct, Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2023), Llama-3.1-70B Grattafiori and et al (2024), Qwen-2-72B-Instruct Qwen et al. (2025), as well as GPT-4o and GPT-4o-mini. Furthermore, we compare CogWriter with two enhanced baselines, SEFL-REFINE Madaan et al. (2023) and Chain-of-Thought prompting Wei et al. (2022), using GPT-4o-mini as the base model. To evaluate the effectiveness of our CogWriter paradigm, we apply it to four representative foundation models: GPT-4o-mini-2024-07-18, GPT-4o-2024-08-06, Qwen-2.5-14B Team (2024), and Llama-3.3-70B Touvron et al. (2024), which represent both closed-source and open-source models.
@@ 7 Related Work
Multi-agent writing has made notable progress in recent years Hu et al. (2025); Pichlmair et al. (2024), showing how agents can collaborate on diverse writing tasks Wang et al. (2024b); Hong et al. (2024). Research has explored heterogeneous agent integrationChen et al. (2025) and educational applicationsShahzad et al. (2024). In academic writing, frameworks like SciAgents Ghafarollahi and Buehler (2024) demonstrate collaboration among specialized agents for complex writing tasks Wang et al. (2024d); D’Arcy et al. (2024); Su et al. (2024), while the Agents’ Room approach Huot et al. (2024) highlights the value of task decomposition in narrative writing. Beyond academic contexts, multi-agent methods have been applied to creative and informational writing, such as Wikipedia-style articles Shao et al. (2024) and poetry Zhang and Eger (2024); Chen et al. (2024). While these methods focus on collaboration, our work applies cognitive writing principles with agents for planning, monitoring, and revisions, enabling flexible adaptation without task-specific training.

### v3

@@ A Cognitive Writing Perspective for Constrained Long-Form
Kaiyang Wan1, Honglin Mu1, Rui Hao2, Haoran Luo3, Tianle Gu1, Xiuying Chen1thanks: Corresponding author., 1MBZUAI, 2University of Chinese Academy of Sciences, 3Nanyang Technological University {Kaiyang.Wan, Xiuying.Chen}@mbzuai.ac.ae
@@ 1 Introduction
LLMs like ChatGPT Achiam et al. (2023) have begun to mirror human-like capabilities across diverse natural language processing tasks Xi et al. (2023); Luo et al. (2024). From crafting concise summaries Chen et al. (2025b, 2024a) to composing structured reports Schmidgall et al. (2025); Wang et al. (2024d), these models can generate coherent text in a single pass Rasheed et al. (2025); Minaee et al. (2024) with a fluency that often rivals human writers. Recent advances have led to models with expanded context windows of up to 128K tokens Pawar et al. (2024), theoretically enabling the generation of extensive documents Bai et al. (2024). However, these models face significant challenges when tasked with generating constrained long-form text under complex constraints, such as following detailed instructions over 10,000 words Wu et al. (2024a). This limitation poses a crucial barrier for applications requiring extended Shi et al. (2024), well-structured content, including creative design proposals, technical documentation, and comprehensive research reports.
@@ 2 A Cognitive Writing Perspective
However, LLMs lack this ability due to their left-to-right, single-pass generation Yao et al. (2023); Wu et al. (2024b). Without the ability to revisit or reorganize previous content, LLMs struggle with global revisions, such as restructuring sections or ensuring consistency across distant parts of the text Bae and Kim (2024); Cheng et al. (2024a, b). This absence of dynamic reviewing often results in long-form outputs with accumulated errors, inconsistencies, or redundant content.
@@ 4 Methodology
Drawing upon our analysis of cognitive writing processes and the identified limitations of single-pass generation approaches, in this section, we propose CogWriter, a training-free framework that equips LLM with cognitive writing capabilities and enables LLMs to tackle complex constrained long-form generation with human-like strategic thinking.
@@ 5.1 Experimental Setup
Experimental Setup
We evaluate our approach across three categories of models and methods. First, we establish baseline performance using several single-pass generation models from the official LongGenBench repository, including LongWriter-Llama3.1-8B Bai et al. (2024), Llama-3.1-8B-Instruct, Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2023), Llama-3.1-70B Grattafiori and et al (2024), Qwen-2-72B-Instruct Qwen et al. (2025), as well as GPT-4o and GPT-4o-mini. Second, we compare against two prominent enhancement methods: SELF-REFINE Madaan et al. (2023) and Chain-of-Thought (CoT) prompting Wei et al. (2022). These methods are applied to four representative foundation models to ensure comprehensive evaluation across different model capabilities and architectures. Finally, to demonstrate the effectiveness of our CogWriter paradigm, we apply it to the same four foundation models: GPT-4o-mini-2024-07-18, GPT-4o-2024-08-06, Qwen-2.5-14B Team (2024), and Llama-3.3-70B Touvron et al. (2024). This selection encompasses closed-source and open-source models with varying parameter scales, enabling us to evaluate CogWriter’s generalizability. For fair comparison, we implement SELF-REFINE and CoT baselines on these same models alongside our proposed framework.
@@ 7 Related Work
Multi-agent writing has made notable progress in recent years Guo et al. (2024); Liu et al. (2024); Song et al. (2024), showing how agents can collaborate on diverse writing tasks Wang et al. (2024b); Hong et al. (2024). Research has explored heterogeneous agent integration Chen et al. (2025a) and educational applications Shahzad et al. (2024). In academic writing, frameworks like SciAgents Ghafarollahi and Buehler (2024) demonstrate collaboration among specialized agents for complex writing tasks Wang et al. (2024d); D’Arcy et al. (2024); Su et al. (2024), while the Agents’ Room approach Huot et al. (2024) highlights the value of task decomposition in narrative writing. Beyond academic contexts, multi-agent methods have been applied to creative and informational writing, such as Wikipedia-style articles Shao et al. (2024) and poetry Zhang and Eger (2024); Chen et al. (2024b). While these methods focus on collaboration, our work applies cognitive writing principles with agents for planning, monitoring, and revisions, enabling flexible adaptation without task-specific training.